# FloydNet: A Learning Paradigm for Global Relational Reasoning

### v1

@@ Abstract
Developing models capable of complex, multi-step reasoning is a central goal in artificial intelligence. While representing problems as graphs is a powerful approach, Graph Neural Networks (GNNs) are fundamentally constrained by their message-passing mechanism, which imposes a local bottleneck that limits global, holistic reasoning. We argue that dynamic programming (DP), which solves problems by iteratively refining a global state, offers a more powerful and suitable learning paradigm. We introduce FloydNet, a new architecture that embodies this principle. In contrast to local message passing, FloydNet maintains a global, all-pairs relationship tensor and learns a generalized DP operator to progressively refine it. This enables the model to develop a task-specific relational calculus, providing a principled framework for capturing long-range dependencies. Theoretically, we prove that FloydNet achieves 3-WL (2-FWL) expressive power, and its generalized form aligns with the k-FWL hierarchy. FloydNet demonstrates state-of-the-art performance across challenging domains: it achieves near-perfect scores (often >99%) on the CLRS-30 algorithmic benchmark, finds exact optimal solutions for the general Traveling Salesman Problem (TSP) at rates significantly exceeding strong heuristics, and empirically matches the 3-WL test on the BREC benchmark. Our results establish this learned, DP-style refinement as a powerful and practical alternative to message passing for high-level graph reasoning.
@@ 1 Introduction
Neural Algorithmic Reasoning (NAR) (Veličković and Blundell, 2021) provides a promising framework for this challenge, aiming to bridge the logic of classical algorithms with the power of deep learning. Graph Neural Networks (GNNs), particularly the Message Passing (MPNN) paradigm (Gilmer et al., 2017), have become a cornerstone of NAR. However, this architectural choice introduces a fundamental bottleneck: local message passing is inherently insufficient for problems demanding global, holistic reasoning. This locality manifests in two key limitations: restricted information flow, where capturing long-range dependencies requires deep GNNs that are prone to over-smoothing and over-squashing (Li et al., 2018; Topping et al., 2022), hindering effective scaling; and limited expressive power, with most MPNNs being bounded by the 1-WL test (Xu et al., 2019). Although graph transformers (GTs) (Dwivedi and Bresson, 2021; Ying et al., 2021) offer global receptive fields, they often rely on complex positional or structural encodings (Ying et al., 2021; Rampášek et al., 2022) to inject inductive bias into the graph, which can be difficult to design and may still underuse the higher-order relational structure of the graph.
@@ 2.1 Initialization of the Pairwise Relationship Representation
$\mathbf{R}^{(0)}_{ij}=\mathrm{MLP}_{\mathrm{init}}\Big(\big[\mathbf{G},\mathbf{X}_{i},\mathbf{X}_{j},\mathbf{E}_{ij}\big]\Big)$ (1)
@@ 2.3 The FloydBlock: A Learned DP Operator
${\bm{R}}^{\prime}_{\vphantom{(l)}}$ $={\bm{R}}^{(l-1)}+\text{PivotalAttention}(\text{Norm}({\bm{R}}^{(l-1)}))$ (2)
${\bm{R}}^{(l)}_{\vphantom{(l)}}$ $={\bm{R}}^{\prime}_{\vphantom{(l)}}+\text{FFN}(\text{Norm}({\bm{R}}^{\prime}_{\vphantom{(l)}}))$ (3)
@@ 2.3.1 Pivotal Attention: The Core Mechanism
$\mathbf{k}_{ijk}$ $=\mathcal{C}(\mathbf{k}_{ij},\mathbf{k}_{jk}),~~~\mathbf{v}_{ijk}=\mathcal{C}(\mathbf{v}_{ij},\mathbf{v}_{jk})$ (4)
@@
For most graph tasks, an additive combine operation is suitable. The multiplicative operation may be useful for some geometric tasks (see Section˜B.1). The overall computational complexity of Pivotal Attention is $\mathcal{O}(N^{3}\cdot d_{r}+N^{2}\cdot d_{r}^{2})$ (see Section˜B.3). A naive implementation of Pivotal Attention would create large intermediate tensors $\mathbf{k}_{ijk}$ and $\mathbf{v}_{ijk}$ of size $\mathcal{O}(N^{3}\cdot d_{r})$, which is memory-intensive for large graphs. To address this, we implemented a dedicated compute kernel that avoids storing these large intermediate tensors, reducing memory usage to $\mathcal{O}(N^{2}\cdot d_{r})$. While this optimization does not alter the theoretical computational cost, it improved training performance by more than 20 times in practice, as hardware memory bandwidth is a limiting factor.
@@ 2.5 General form of FloydNet (k-FloydNet)
FloydNet is the special case where k = 2, and self-attention of the Transformer is the special case where k=1. The expressive power of k-FloydNet increases with the value of $k$, as higher-order interactions among sets of vertices can be captured more effectively. The overall computational complexity of Pivotal Attention is $\mathcal{O}(N^{k+1}\cdot d_{r}+N^{k}\cdot d_{r}^{2})$ (see Section˜B.2 and Section˜B.3).
@@ 4.2 Neural Algorithmic Reasoning
On held-out general TSP instances with $100<N\leq 200$, our model demonstrates a strong ability to significantly improve the optimality rate through multiple sampling as shown in Figure˜4(a). Specifically, it increases the optimality rate from $83.6\%$ to $99.8\%$ after 10 samples, substantially surpassing the $38.8\%$ achieved by the strong Linkern heuristic. Furthermore, it exhibits excellent extrapolation capabilities concerning the number of nodes ($N$). As $N$ increases, the optimality rate of Linkern rapidly deteriorates, dropping to $16.1\%$ for instances where $180\leq N\leq 200$. In sharp contrast, our model with 10 samples maintains a robust optimality rate of $99.4\%$ in this challenging range. We also examined a subset containing only single-solution instances. On non metric TSP instances with $100<N\leq 200$, the FloydNet finds the exact optimal tour in $92.6\%$ of cases, as shown in Figure˜4(b). This performance markedly outperforms the Linkern heuristic, which achieves only a $15.7\%$ optimality rate on these larger instances. The performance gap widens significantly with instance size: Linkern’s optimality rate plummets to $1.3\%$ when $180\leq N\leq 200$, while FloydNet still achieves $88.3\%$. For smaller validation instances ($N\leq 100$), FloydNet demonstrates robust performance within the training distribution, with an optimality rate exceeding $96\%$. On Metric TSP instances, the Linkern heuristic performs substantially better than on Non-Metric TSP instances, leveraging the underlying geometric properties. In this setting, FloydNet achieves an optimality rate that is comparable to that of the Linkern heuristic.
@@
We analyzed the failure cases of FloydNet on singe-solution instances and found that only a small fraction ($0.8\%$) corresponded to valid but suboptimal TSP tours, $6.2\%$ produced multiple disjoint cycles covering all nodes, and the remaining $93\%$ generated fewer edges than nodes, resulting in incomplete tours that failed to visit all nodes. We hypothesize that this phenomenon may stem from numerical issues or interactions with mechanisms such as initialization, normalization, or optimization. Notably, scaling up the model effectively mitigates the effect, highlighting a promising direction for future study.
@@ 5 Related Work
5.1 Architectures Beyond 1-WL Expressiveness
A central theme in GNN research is overcoming the expressive limitations of MPNNs, famously bounded by the 1-WL test (Xu et al., 2019; Morris et al., 2019). This pursuit has led to a variety of more powerful architectures, including Subgraph GNNs (Bevilacqua et al., 2022; Zhang and Li, 2021), Spectral GNNs(Zhang et al., 2024b; Lim et al., 2023) and higher-order GNNs(Maron et al., 2019b; a) grounded in more powerful theoretical tests. Among these, the k-Folklore Weisfeiler-Lehman (k-FWL) test (Cai et al., 1992; Maron et al., 2019a) serves as a key benchmark for models that reason over tuples of nodes. The 2-FWL test, for instance, refines representations for pairs of nodes $(u,v)$ by aggregating over all intermediate pivot nodes $w$, considering relational triangles of the form $u\leadsto w\leadsto v$. Architectures like the Folklore GNN (FGNN) (Zhang et al., 2024a) are designed to mimic this process.
While many expressive models rely on enriching local structures or implementing complex tensor operations, FloydNet introduces a distinct computational paradigm. It is based on learning a global, DP-style operator that is algorithmically aligned with the pivotal aggregation pattern of the FWL tests. Our empirical results strongly support a key theoretical connection: the relational refinement in k-FloydNet is not merely a novel heuristic but is empirically equivalent in expressive power to the k-FWL test. We give a formal discussion on Appendix˜C. This result precisely characterizes k-FloydNet, positioning it as an elegant and practical realization of this powerful theoretical class.
5.2 Transformers and Positional Encodings
The standard Transformer architecture (Vaswani et al., 2017) excels at processing 1D sequences. However, graphs lack a canonical ordering, making the direct application of Transformers challenging. Graph Transformers (GTs) (Dwivedi and Bresson, 2021; Ying et al., 2021; Kim et al., 2022; Rampasek et al., 2022; Lai et al., 2023) adapt this architecture by treating nodes as an unordered set, but this makes them inherently blind to the graph’s topology. To compensate, they require carefully designed Positional and Structural Encodings (PSEs) to re-inject the structural information that was lost.
This work proposes a more principled generalization. Instead of operating on a 1D set of node representations, FloydNet’s Pivotal Attention mechanism operates directly on the 2D relational space of all node pairs. This perspective sidesteps the need for PSEs entirely. While many PSEs have been developed, often based on Laplacian eigenvectors (Kreuzer et al., 2021), they introduce their own challenges, such as eigenvector ambiguity (Lim et al., 2023). Even more robust spectral features (Huang et al., 2024; Li et al., 2020) have fundamental theoretical limits; a comprehensive analysis by Zhang et al. (2024b) shows their expressive power is strictly weaker than the 3-WL test. Thus, most GTs, despite their architectural complexity, operate below the expressiveness ceiling that FloydNet empirically achieves. By building the graph’s relational structure into its core computational block, FloydNet provides an elegant and powerful alternative that is naturally permutation-equivariant.
@@ 6 Conclusion and Future Work
This paper introduced FloydNet, a new paradigm for graph reasoning based on a learned, DP-style global refinement operator that replaces local message passing. Theoretically, FloydNet and its’ extends (k-FloydNet) is equivalent to the powerful 2-FWL(k-FWL) test, yet empirically, it achieves such distinguishing power on challenging benchmarks. This expressiveness translates to state-of-the-art performance on algorithmic reasoning (CLRS-30) and combinatorial optimization (TSP), establishing our approach as a powerful and practical paradigm for global graph reasoning.
Future work can focus on improving scalability beyond the current $\mathcal{O}(N^{3})$ complexity through sparse mechanisms, extending the framework to new domains like multimodality for which its relational structure is a natural fit, and further strengthening its theoretical foundations by connecting to Abstract DP theory(Bertsekas, 2022; 2025). Ultimately, we envision FloydNet as a core component of future System-2 reasoning systems.
@@ A.2 Dataset Details
The CLRS-30 benchmark (Veličković et al., 2022) is a comprehensive suite of algorithmic reasoning tasks, serving as a standardised testbed for evaluating a model’s ability to learn classical algorithms. It is designed to directly assess a model’s capacity for complex reasoning and out-of-distribution (OOD) generalization, where OOD specifically refers to executing a learned algorithm on inputs of a larger size (e.g., training on arrays of size 16 and testing on size 64). The benchmark is quantitatively composed of 30 classical algorithmsrom the textbook introduction to Algorithms(Cormen et al., 2022), which are grouped into 8 distinct categories such as sorting, searching, and dynamic programming. The core logical format of the dataset represents every algorithm’s execution as a trajectory on a graph. Each data sample contains the problem’s initial inputs, the final correct outputs, and a time-series of intermediate states called hints. Crucially, all these components are mapped to a unified graph representation as node, edge, or global graph features. This standardized format is the key that enables a single GNN processor to be trained across multiple, structurally diverse algorithms, a central goal of our work.
@@ B.1 Multiplicative Pivotal Attention
subsectionLimitations and Engineering Optimization The primary trade-off of the FloydNet architecture is its computational complexity. The global, all-pairs refinement via Pivotal Attention scales as $\mathcal{O}(N^{3})$, which is inherent to the 3-WL expressive power it achieves. While this prohibits direct application to massive graphs (e.g., millions of nodes), it is well-suited for high-value reasoning tasks on dense graphs of moderate size (e.g., algorithmic reasoning, molecular modeling).
@@
B.2 Details of the Higher-Order K-FloydNet Framework
@@
For example, to initialize the tensor for 3-FloydNet using graph feature $G$, node feature $X$, edge feature $E$ and 3-dimensional hyper-edge feature $H$ (which may be unavailable), we can concatenate all ${3\choose 0}+{3\choose 1}+{3\choose 2}+{3\choose 3}=8$ possible lower-order relationships within the triplet $(i,j,k)$:
@@ B.3 Complexity of FloydNet
B.4 Comparison with Transformers and the Canonical Ordering Problem

### v2

@@ Abstract
Developing models capable of complex, multi-step reasoning is a central goal in artificial intelligence. While representing problems as graphs is a powerful approach, Graph Neural Networks (GNNs) are fundamentally constrained by their message-passing mechanism, which imposes a local bottleneck that limits global, holistic reasoning. We argue that dynamic programming (DP), which solves problems by iteratively refining a global state, offers a more powerful and suitable learning paradigm. We introduce FloydNet, a new architecture that embodies this principle. In contrast to local message passing, FloydNet maintains a global, all-pairs relationship tensor and learns a generalized DP operator to progressively refine it. This enables the model to develop a task-specific relational calculus, providing a principled framework for capturing long-range dependencies. Theoretically, we prove that FloydNet achieves 3-WL expressive power (equivalently, 2-FWL), and that its higher-order generalization aligns with the k-FWL hierarchy. FloydNet demonstrates state-of-the-art performance across challenging domains: it achieves near-perfect scores (often >99%) on the CLRS-30 algorithmic benchmark, finds exact optimal solutions for the general Traveling Salesman Problem (TSP) at rates significantly exceeding strong heuristics, and empirically matches the 3-WL test on the BREC benchmark. Our results establish this learned, DP-style refinement as a powerful and practical alternative to message passing for high-level graph reasoning.
@@ 1 Introduction
Neural Algorithmic Reasoning (NAR) (Veličković and Blundell, 2021) provides a promising framework for this challenge, aiming to bridge the logic of classical algorithms with the power of deep learning. Graph Neural Networks (GNNs), particularly the Message Passing (MPNN) paradigm (Gilmer et al., 2017), have become a cornerstone of NAR. However, this architectural choice introduces a fundamental bottleneck: local message passing is inherently insufficient for problems demanding global, holistic reasoning. This locality manifests in two key limitations: restricted information flow, where capturing long-range dependencies requires deep GNNs that are prone to over-smoothing and over-squashing (Li et al., 2018; Topping et al., 2022), hindering effective scaling; and limited expressive power, with most MPNNs being bounded by the 1-WL test (Xu et al., 2019). Although graph transformers (GTs) (Dwivedi and Bresson, 2021; Ying et al., 2021) offer global receptive fields, they often rely on complex positional or structural encodings (Rampášek et al., 2022) to inject inductive bias into the graph, which can be difficult to design and may still underuse the higher-order relational structure of the graph.
@@ 2.1 Initialization of the Pairwise Relationship Representation
${\bm{R}}^{(0)}_{ij}=\mathrm{MLP}_{\mathrm{init}}\Big(\big[\mathbf{G},\mathbf{X}_{i},\mathbf{X}_{j},\mathbf{E}_{ij}\big]\Big)$ (1)
@@ 2.3 The FloydBlock: A Learned DP Operator
${\bm{R}}^{\prime}$ $={\bm{R}}^{(l-1)}+\text{PivotalAttention}(\text{Norm}({\bm{R}}^{(l-1)}))$ (2)
${\bm{R}}^{(l)}$ $={\bm{R}}^{\prime}+\text{FFN}(\text{Norm}({\bm{R}}^{\prime}))$ (3)
@@ 2.3.1 Pivotal Attention: The Core Mechanism
$\mathbf{k}_{ijk}$ $=\mathcal{C}(\mathbf{k}_{ij},\mathbf{k}_{jk}),$ $\mathbf{v}_{ijk}$ $=\mathcal{C}(\mathbf{v}_{ij},\mathbf{v}_{jk})$ (4)
@@
For most graph tasks, an additive combine operation is suitable. The multiplicative operation may be useful for some geometric tasks (see Section˜B.1). The overall computational complexity of Pivotal Attention is $\mathcal{O}(N^{3}\cdot d_{r}+N^{2}\cdot d_{r}^{2})$ (see Section˜B.4). A naive implementation of Pivotal Attention would create large intermediate tensors $\mathbf{k}_{ijk}$ and $\mathbf{v}_{ijk}$ of size $\mathcal{O}(N^{3}\cdot d_{r})$, which is memory-intensive for large graphs. To address this, we implemented a dedicated compute kernel that avoids storing these large intermediate tensors, reducing memory usage to $\mathcal{O}(N^{2}\cdot d_{r})$. While this optimization does not alter the theoretical computational cost, it improved training performance by more than 20 times in practice, as hardware memory bandwidth is a limiting factor.
@@ 2.5 General form of FloydNet (k-FloydNet)
FloydNet is the special case where k = 2, and self-attention of the Transformer is the special case where k=1. The expressive power of k-FloydNet increases with the value of $k$, as higher-order interactions among sets of vertices can be captured more effectively. The overall computational complexity of Pivotal Attention is $\mathcal{O}(N^{k+1}\cdot d_{r}+N^{k}\cdot d_{r}^{2})$ (see Section˜B.3 and Section˜B.4).
@@ 4.2 Neural Algorithmic Reasoning
On held-out general TSP instances with $100<N\leq 200$, our model demonstrates a strong ability to significantly improve the optimality rate through multiple sampling as shown in Figure˜4(a). Specifically, it increases the optimality rate from $83.6\%$ to $99.8\%$ after 10 samples, substantially surpassing the $38.8\%$ achieved by the strong Linkern heuristic. Furthermore, it exhibits excellent extrapolation capabilities concerning the number of nodes ($N$). As $N$ increases, the optimality rate of Linkern rapidly deteriorates, dropping to $16.1\%$ for instances where $180\leq N\leq 200$. In sharp contrast, our model with 10 samples maintains a robust optimality rate of $99.4\%$ in this challenging range. We also examined a subset containing only single-solution instances. On non-metric TSP instances with $100<N\leq 200$, the FloydNet finds the exact optimal tour in $92.6\%$ of cases, as shown in Figure˜4(b). This performance markedly outperforms the Linkern heuristic, which achieves only a $15.7\%$ optimality rate on these larger instances. The performance gap widens significantly with instance size: Linkern’s optimality rate plummets to $1.3\%$ when $180\leq N\leq 200$, while FloydNet still achieves $88.3\%$. For smaller validation instances ($N\leq 100$), FloydNet demonstrates robust performance within the training distribution, with an optimality rate exceeding $96\%$. On Metric TSP instances, the Linkern heuristic performs substantially better than on Non-Metric TSP instances, leveraging the underlying geometric properties. In this setting, FloydNet achieves an optimality rate that is comparable to that of the Linkern heuristic.
@@
We analyzed the failure cases of FloydNet on single-solution instances and found that only a small fraction ($0.8\%$) corresponded to valid but suboptimal TSP tours, $6.2\%$ produced multiple disjoint cycles covering all nodes, and the remaining $93\%$ generated fewer edges than nodes, resulting in incomplete tours that failed to visit all nodes. We hypothesize that this phenomenon may stem from numerical issues or interactions with mechanisms such as initialization, normalization, or optimization. Notably, scaling up the model effectively mitigates the effect, highlighting a promising direction for future study.
@@ 5 Related Work
5.1 Beyond 1-WL: the WL Hierarchy and Higher-Order GNNs
A central goal in GNN research is to overcome the expressive limits of message passing, which are bounded by the 1-WL test (Xu et al., 2019; Morris et al., 2019). Representative directions include higher-order networks (Maron et al., 2019b; a), subgraph-based models (Bevilacqua et al., 2022; Zhang and Li, 2021), and spectral approaches (Zhang et al., 2024b; Lim et al., 2023). The k-Folklore Weisfeiler–Lehman (k-FWL) hierarchy (Cai et al., 1992; Maron et al., 2019a) provides a unifying yardstick: it refines k-tuple colors by aggregating over pivot-generated neighbor tuples; FGNN (Zhang et al., 2024a) is an example explicitly designed to mimic this refinement.
Our work follows an algorithmically aligned perspective: k-FloydNet learns a global DP-style refinement operator whose core pivot aggregation matches k-FWL. Empirically, k-FloydNet matches k-FWL expressivity (see Appendix˜C), providing a clean characterization of the architecture class.
5.2 1D/2D/kD Transformer-Style Models
We view many Transformer-style graph models as operating over different-dimensional state spaces. 1D (node-set attention). Graph Transformers perform attention over a set of node states (Dwivedi and Bresson, 2021; Ying et al., 2021; Kim et al., 2022; Rampasek et al., 2022; Lai et al., 2023) and typically require positional/structural encodings to re-introduce topology.
2D (pair/edge attention). Edge Transformer (Bergen et al., 2021) lifts the state space to node pairs and applies triangular (2-hop) updates through pivots; Müller et al. (2024) connects its expressivity to the WL hierarchy. Related approaches such as EGT introduce an explicit residual edge channel alongside node attention (Hussain et al., 2021).
Our k-FloydNet unifies these views: $k=1$ recovers the 1D Transformer perspective, $k=2$ captures the 2D pairwise paradigm while making the “Join-V vs. Join-K” distinction explicit (Section˜B.5.1), and $k>2$ extends to higher-order relational reasoning with a k-FWL characterization.
@@ 6 Conclusion and Future Work
This paper introduced FloydNet, a new paradigm for graph reasoning based on a learned, DP-style global refinement operator that replaces local message passing. Theoretically, FloydNet (i.e., 2-FloydNet) matches 2-FWL expressivity (equivalently, 3-WL), and its higher-order generalization k-FloydNet aligns with the k-FWL hierarchy. Empirically, this expressiveness translates into strong performance on algorithmic reasoning (CLRS-30) and combinatorial optimization (TSP), establishing our approach as a practical paradigm for global graph reasoning.
Future work includes improving scalability beyond the current $\mathcal{O}(N^{3})$ complexity via sparse mechanisms, extending the framework to new domains such as multimodality where relational structure is central, and strengthening theoretical foundations by connecting to abstract dynamic programming (Bertsekas, 2022; 2025). Ultimately, we envision FloydNet as a core component of future System-2 reasoning systems.
@@ A.2 Dataset Details
The CLRS-30 benchmark (Veličković et al., 2022) is a comprehensive suite of algorithmic reasoning tasks, serving as a standardised testbed for evaluating a model’s ability to learn classical algorithms. It is designed to directly assess a model’s capacity for complex reasoning and out-of-distribution (OOD) generalization, where OOD specifically refers to executing a learned algorithm on inputs of a larger size (e.g., training on arrays of size 16 and testing on size 64). The benchmark is quantitatively composed of 30 classical algorithms from the textbook introduction to Algorithms(Cormen et al., 2022), which are grouped into 8 distinct categories such as sorting, searching, and dynamic programming. The core logical format of the dataset represents every algorithm’s execution as a trajectory on a graph. Each data sample contains the problem’s initial inputs, the final correct outputs, and a time-series of intermediate states called hints. Crucially, all these components are mapped to a unified graph representation as node, edge, or global graph features. This standardized format is the key that enables a single GNN processor to be trained across multiple, structurally diverse algorithms, a central goal of our work.
@@ B.1 Multiplicative Pivotal Attention
B.2 Limitations and Engineering Optimization
The primary trade-off of the FloydNet architecture is its computational complexity. The global, all-pairs refinement via Pivotal Attention scales as $\mathcal{O}(N^{3})$, which is inherent to the 3-WL expressive power it achieves. While this prohibits direct application to massive graphs (e.g., millions of nodes), it is well-suited for high-value reasoning tasks on dense graphs of moderate size (e.g., algorithmic reasoning, molecular modeling).
@@ B.2 Limitations and Engineering Optimization
B.3 Details of the Higher-Order K-FloydNet Framework
@@
For example, to initialize the tensor for 3-FloydNet using graph feature $G$, node feature $X$, edge feature $E$ and 3-dimensional hyper-edge feature $H$ (which may be unavailable), we can concatenate all $\binom{3}{0}+\binom{3}{1}+\binom{3}{2}+\binom{3}{3}=8$ possible lower-order relationships within the triplet $(i,j,k)$:
@@ B.4 Complexity of FloydNet
B.5 Comparison with Transformers and the Canonical Ordering Problem
B.5.1 Relationship to Edge Transformer: Join-V vs. Join-K
Edge Transformer (Bergen et al., 2021) is based on triangular attention, which can be viewed as a Join-V aggregation pattern: for a target pair $(i,k)$ and pivot $j$, the update joins value signals along the triangle (e.g., $V=\{\mathbf{v}_{ij},\mathbf{v}_{jk},\mathbf{v}_{ik}\}$). In contrast, FloydNet uses Pivotal Attention with a Join-K pattern, where the pivot-conditioned compatibility is constructed by joining keys (e.g., $K=\{\mathbf{k}_{ij},\mathbf{k}_{jk},\mathbf{k}_{ik}\}$) before performing attention.
This distinction aligns with our theory: to simulate k-FWL, the model must form pivot-conditioned keys that index neighbor tuples (a “Join-K” pattern), which motivates the design of Pivotal Attention.